# Training LLMs over Neurally Compressed Text

### v1

@@ 1 Introduction
Today’s large language models (LLMs) are almost exclusively trained over subword tokens. The tokenizers used to produce these tokens—often BPE [23, 56] or Unigram [37], as implemented by the SentencePiece library [38]—are compressors that typically achieve ~4$\times$ compression over natural language text [74]. [^1] We refer here to “token-level” compression rate, i.e., the length reduction between a raw UTF-8 byte sequence and the corresponding sequence of subword tokens. If instead we measure the number of bits required to encode the two sequences, subword compression typically delivers ~2$\times$ or less compression, depending on vocabulary size, which typically ranges from 32k to 256k. See Section 3.4 for discussion. While these tokenizers “hide” the character-level makeup of each token from the LLM [74, 44], this downside is widely seen as outweighed by the significant benefits of compression. Compared to raw byte-level models, an LLM trained over subword tokens sees ~4$\times$ more text per token, allowing it to model longer-distance dependencies, ingest more pretraining data, and predict more text at inference time, all without increasing compute. [^2] The increased cost of the input embedding and final softmax layers due to increased vocabulary size is negligible for all but the smallest models.
Given these advantages, it raises the question, could we compress text further to achieve even greater gains? It is well known that autoregressive language models can be turned into lossless text compressors, and recent work has shown that LLMs can easily achieve 12$\times$ compression over English text [16]. [^3] Specifically, the authors show that Chincilla 70B [30] can compress 2048-byte subspans of enwik9 at a 12$\times$ bit-level compression rate. Can we simply train an LLM over this neurally compressed text?
@@
In this paper we explore various options for doing so, focusing primarily on the idea of using Arithmetic Coding (AC) [73], which is known to reach the near-optimal compression rate for a particular model that assigns probabilities to text continuations. Figure 1 presents our high-level approach. First, a small language model “M1” is trained over raw byte sequences. Next, this frozen model is used to compress pretraining corpus text by applying a standard compression algorithm like AC. The resulting compressed bitstream is then chunked into tokens, which are used to train “M2”, a language model that directly reads and writes neural-compressed text.
Given a perfect probabilistic model of the raw byte sequence, the compression step would output a fully-compressed bitstream that would be indistinguishable from random noise, and hence unlearnable by M2. In reality, M1 can never be perfect [78], so the M1-compressed output will still contain learnable patterns. We explore whether using compression powered by a relatively small M1 is able to “remove” the simple structure that M1 understands from the input—e.g., patterns of spelling, word frequency, and basic grammar—while retaining any higher-level structure that M1 fails to model—e.g., patterns requiring “deeper” reasoning and long range coherence. A larger M2 would then learn to model this higher-level structure, without needing to relearn the low-level structure removed by M1.444Intuitively, training M2 could be seen as analogous to fitting the residuals of M1 [21]. In theory, this process could be repeated by training an even-larger M3 model on text compressed by M2, and so on.
@@
At the same time, our best M2 models underperform subword baselines. We suspect this is due at least in part to the relatively unstable mappings our neural tokenizers induce between words and tokens. By contrast, standard subword tokenizers induce essentially stable word-to-token mappings, which likely makes the token sequences they output well-suited for LLM training. [^5] See Appendix L for some counterexamples to subword tokenizers producing stable word-to-token mappings. We illustrate this contrast through qualitative examples. Whether a neural tokenizer can reach a high level of compression while maintaining high learnability for LLM training is an interesting question for future research.
Our main contributions are as follows: (1) Outline advantages and challenges of training over neural compressed text. (2) Compare LLMs trained over different tokenizers along two axes: bits/byte and FLOPs/byte. (3) Show that standard LLMs can’t learn to model vanilla AC-compressed text. (4) Show that GZip-compressed text is learnable by standard LLMs, but not competitive. (5) Propose compression via Equal-Info Windows, and show that it enables learning over neural compressed text.
@@ 2 Motivation and Background
2.1 Advantages of Training over Neural-Compressed Text
@@
The most straightforward advantage is efficiency. By compressing the same text into a shorter token sequence, the model can process more text for the same computational cost. In particular, a model trained over $C{\mkern-1.0mu\times\mkern-1.0mu}{}$ compressed text will see $C{\mkern-1.0mu\times\mkern-1.0mu}{}$ more text during training compared to a model trained over raw text, given an equal compute budget. Increasing the amount of data seen in pretraining is often an effective means of improving performance [35, 30]. Processing text more efficiently also confers benefits at inference time, reducing the serving cost for handling a request of a given prompt and continuation length. In addition to reducing the raw compute needed for inference, compression can also improve inference latency, since generating better-compressed output requires fewer sequential autoregressive steps.
@@
A second advantage is that working with compressed text allows modeling longer contextual dependencies. In vanilla transformer-based models, computation for the self-attention layer scales quadratically with the sequence length, $O(n^{2}d)$. This has limited the sequence lengths used by such models in practical settings to ~$10$k tokens. [^6] Exploring sub-quadratic attention mechanisms is an area of active research [1, 70, 36, 75, 5, 9, et alia]. However, regardless of the cost of attention, compressing the input increases the effective context “for free”. If, via compression, each token represents (on average) $C$ bytes of raw text, then the resulting LLM can model dependencies across $C{\mkern-1.0mu\times\mkern-1.0mu}{}$ longer distances compared to a raw text model operating over the same token sequence length. While the benefits of modeling longer context (beyond ~$1{,}000$ bytes) are modest when viewed merely as perplexity gains [51], the ability to condition on long context is critical for many applications, such as retrieving content from a document, or answering a coding question provided documentation.
@@
A third potential advantage of training over compressed text is that information will be spread more uniformly across the sequence. By the nature of compression, a text span that is relatively predictable (e.g., a boilerplate notice) will be more compressible than a span with high perplexity (e.g., a unique product serial number). When an LLM is trained over well-compressed text, each token will represent roughly an equal amount of information. Since the LLM allocates equal compute to each token, this amounts to allocating more compute for “harder” text spans. This adaptivity is similar in spirit to “Adaptive Computation Time” (ACT) [27], which learns to allocate additional compute at some sequence positions in an end-to-end manner, but with the advantage that in our case the computation remains “dense”—identical operations are applied at each position. [^7] It should be noted that ACT learns to allocate more compute where it is useful, as opposed to merely where the predictions are hard. For example, ACT learns to not waste compute on inherently unpredictable text spans. We expect that as a heuristic, allocating more compute to higher-perplexity text spans is valuable, but leave this to future work to verify.
@@ 2.2 Challenges of Training over Compressed Text
As a prerequisite for M2 to effectively predict continuations of compressed text, we anticipate that it is necessary for M2 to have the ability to decompress bits $\rightarrow$ text and compress text $\rightarrow$ bits. These sub-tasks are challenging in their own right. First, M2 needs to accurately “simulate” M1 in order to know the probabilities it assigns to the text, which determine the output of compression. [^8] For Arithmetic Coding, not only would M2 need to know the probabilities M1 assigns to the observed text, but it would also need to know the probabilities assigned to many unobserved symbols. This is because Arithmetic Coding operates over cumulative probabilities, i.e., the probability that the next symbol is e or any alphabetically preceding symbol. Training models to mimic other models can be difficult [41], and even in settings where models do learn to copy the behavior of another network [29], this is often only when looking at which symbol was assigned the highest probability—the actual probabilities assigned often differ [60]. Second, M2 needs to learn the compression procedure itself. In our case, this means tracking the Arithmetic Coding algorithm, which requires maintaining high-precision numerical state across long contexts. We investigate these sub-tasks in detail in Section 5.2.
A further learnability challenge is the high level of context sensitivity needed to interpret a bitstream of compressed text. When chunked into tokens, a particular bit subsequence (e.g., 10111001) can map onto the same token despite having no stable “meaning” across occurrences. We show examples in Section 6.1, where a token maps to many different underlying text forms, necessitating strong contextual understanding. While LLMs are robust to some level of polysemy, as highlighted by the success of Hash Embeddings [62] where multiple unrelated words share a single token representation, we suspect this has its limits.
@@
Finally, a specific challenge of training over neural compressed text is that multiple models need to be stored and run side-by-side in order to perform inference. We assume that if M1 is relatively small, this additional overhead is not a significant drawback compared to a standard tokenizer, which is also a separate model that is needed to tokenize text input and detokenize LLM outputs. In evaluating our approach, we include M1 compute in our calculations of total inference cost (FLOPs/byte).
@@ 2.3 Compression
In this work, we focus on lossless compression, which aims to encode a sequence of input symbols, $x_{0:N}=\{x_{0},x_{1},\dots,x_{N}\}\in X^{|V|}$, into a bitstream while minimizing the expected length of the bitstream. Compression methods are often factored into a “modeling” component and a “coding” component [45]. The input sequence can be viewed as a sample from a true distribution $p$, $x_{0:N}\sim\leavevmode\nobreak\ p$, with a standard autoregressive decomposition, $p(x_{0:N})=\prod_{i=1}^{N}p(x_{i}|x_{0},\dots,x_{i-1})$. The “modeling” component aims to approximate $p$ with $\hat{p}$. While some compression algorithms assume static probabilities for each symbol, stronger algorithms are “adaptive”, meaning that symbol probabilities may change based on context. In this work, we use context-aware transformer-based language models to represent $\hat{p}$.
The “coding” component of a compression algorithm converts the input sequence to a bitstream of length $\ell(x_{0:N})$. To maximize compression, we want a coding algorithm that minimizes the expected number of bits in the bitstream, $L\coloneqq\mathbb{E}_{x_{0:N}\sim p}[\ell(x_{0:N})]$. This is done by assigning shorter bit sequences to common symbols and longer sequences to less common ones. [^9] This process can result in extremely uncommon sequences becoming longer under compression, as no algorithm can compress all possible input strings [45]. In practice, natural language inputs are highly compressible and these edge cases are inputs that one would not recognize as natural language. The expected length is lower bounded by $L\geq H(p)$ where $H(p)\coloneqq\mathbb{E}_{x_{0:N}\sim p}[-\log_{2}p(x)]$ [57]. This means that, given a near-optimal coding algorithm, the achievable level of compression derives from how well the model $\hat{p}$ approximates $p$.
@@ 2.4 Arithmetic Coding
Arithmetic Coding [53, 49] uses a model $\hat{p}$ to compresses a sequence $x_{0:N}$ to a bitstream, which is the binary expansion of a float $f\in[0,1)$. The float $f$ is found by assigning successively smaller sub-intervals to each symbol $x_{i}\in x_{0:N}$, with the final interval enclosing $f$. An interval is made of an upper and lower bound, $I_{i}=[l_{i},u_{i})$ and its size is given by $u_{i}-l_{i}$. Starting with $I_{0}=[0,1)$, at each step of encoding, the interval for the symbol $x_{i}$ is created by partitioning the interval $I_{i-1}$ based on the cumulative distribution of $\hat{p}$ given the previous context, $\hat{p}_{cdf}(x_{i}|x_{<i})$. The size of this interval is given by $\text{size}(I_{i-1})*\hat{p}(x_{i}|x_{<i})$. Thus:
@@
The coding component of Arithmetic Coding is nearly optimal: the output bitstream will have a length of $-\lceil\log\hat{p}(x_{0:N})\rceil+1$ bits when using infinite precision. In the finite precision setting using $\beta$ bits, an extra $O(N2^{-\beta})$ bits are added [31]. See [73] for an example implementation. In our experiments, we use precision $\beta=14$. The practical effect of using a finite precision implementation of Arithmetic Coding is that the model’s cumulative distribution gets quantized to integers using $\beta$ bits. This results in a minimum probability of $2^{-\beta}$ being assigned to all tokens.
@@ 2.5 Related Work
Recent work has looked at using large language models for compression, but has not to our knowledge attempted to train subsequent models over the resulting compressed output. Works like [16] use a transformer language model as the modeling component of Arithmetic Coding, but they do not train over compressed output nor do they make modifications to the compression algorithm to facilitate learnability by downstream models. Additionally, they focus on the setting of compressing fixed-size sequences of bytes. By contrast, our models operate over input sequences of fixed token length. This allows for models with higher compression rates to leverage longer contexts, as more bytes are included in the input.
@@
[63] proposes changes to Arithmetic Coding to make it more amenable to use with LLMs—namely, they rank sort the logits from the model before creating text intervals, $I_{i}(x_{0:N})$. This could help alleviate issues stemming from errors in M2’s simulation of M1. However, they do not train models on top of their compressed output.
@@
Some approaches to “token-free” (i.e., purely character- or byte-level) language modeling down-sample the input sequence via convolutions [13, 61], which could be seen as a form of end-to-end neural tokenization. However one important distinction is that the resulting tokenization is “soft”—outputting high-dimensional vectors and not implying a discrete segmentation—in contrast to our tokenization that outputs discrete tokens.
@@
Methods for learning discrete tokenization end-to-end have also been proposed [11, 25]. In the case of MANTa [25], the learned segmentation appears to be fairly semantic (i.e., respecting word and morpheme boundaries), which could be an advantage over our approach. However, they lack our bias towards encoding an equal amount of information per token.
@@
In modeling audio, it is common practice to use learned tokenizers that compress the raw input signal to discrete tokens from a fixed-size codebook [64, 3, 12, 6]. However, this compression is lossy, whereas we focus on lossless compression.
@@
Other recent work focuses on using the “modeling” component from well-known compressors to do other tasks. [34] uses the model from GZip to perform text classification. [68] uses the Arithmetic Decoding algorithm with an LLM as the model to do diverse parallel sampling from that LLM. One could imagine that the “model” of our compressors (M1) is a teacher for M2, but unlike these other applications, the M1 values are not used outside of compression.
@@
[40] also explores learning over compressed text, but with several key differences. First, they use n-gram language models [57] while we use LLMs. Second, their model is conditioned on compressed bitstreams but produces a distribution over the raw, uncompressed, bytes while our M2 models predict directly in the compressed space. Additionally, they only consider static Huffman coding [32] as the algorithm to compress model inputs. While this avoids the context sensitivity issues we outline in Section 2.2, it results in a far worse compression rate compared to the adaptive compression methods we use. One important distinction is that their equal-information windows are overlapping, and used as a sliding window to provide context to their n-gram language model. By contrast our equal-information windows are non-overlapping, and used to segment text into a series of equal-length bitstrings that can be interpreted independently by M2, and whose boundaries are easily identifiable, as they map to a fixed number of M2 tokens.
@@
Concurrently, [26] explores how the compression performance of a tokenizer correlates with downstream model performance. They find that tokenizers that compress better perform better, which generally aligns with our findings, particularly in the large vocabulary setting, see Fig. 6. However, we find that using the strongest compressors is detrimental to learnability, as seen in the AC line in Fig. 3. These conflicting results likely stem from differences in tokenization strategy. Their work is restricted to BPE-based compressors while we explore stronger compressors built on LLMs and Arithmetic Coding. The qualitative differences between these classes of tokenizers are explored more in Section 6.1.

### v2

@@ 1 Introduction
Today’s large language models (LLMs) are almost exclusively trained over subword tokens. The tokenizers used to produce these tokens—often BPE (Gage, 1994, Sennrich et al., 2016) or Unigram (Kudo, 2018), as implemented by the SentencePiece library (Kudo & Richardson, 2018)—are compressors that typically achieve ~4$\times$ compression over natural language text (Xue et al., 2022). [^1] We refer here to “token-level” compression rate, i.e., the length reduction between a raw UTF-8 byte sequence and the corresponding sequence of subword tokens. If instead we measure the number of bits required to encode the two sequences, subword compression typically delivers ~2$\times$ or less compression, depending on vocabulary size, which typically ranges from 32k to 256k. See Section 3.4 for discussion. While these tokenizers “hide” the character-level makeup of each token from the LLM (Xue et al., 2022, Liu et al., 2023), this downside is widely seen as outweighed by the significant benefits of compression. Compared to raw byte-level models, an LLM trained over subword tokens sees ~4$\times$ more text per token, allowing it to model longer-distance dependencies, ingest more pretraining data, and predict more text at inference time, all without increasing compute. [^2] The increased cost of the input embedding and final softmax layers due to increased vocabulary size is negligible for all but the smallest models.
Given these advantages, it raises the question, could we compress text further to achieve even greater gains? It is well known that autoregressive language models can be turned into lossless text compressors, and recent work has shown that LLMs can easily achieve 12$\times$ compression over English text (Delétang et al., 2024). [^3] Specifically, the authors show that Chincilla 70B (Hoffmann et al., 2022) can compress 2048-byte subspans of enwik9 at a 12$\times$ bit-level compression rate. Can we simply train an LLM over this neurally compressed text?
@@
In this paper we explore various options for doing so, focusing primarily on the idea of using Arithmetic Coding (AC) (Witten et al., 1987), which is known to reach the near-optimal compression rate for a particular model that assigns probabilities to text continuations. Figure 1 presents our high-level approach. First, a small language model “M1” is trained over raw byte sequences. Next, this frozen model is used to compress pretraining corpus text by applying a standard compression algorithm like AC. The resulting compressed bitstream is then chunked into tokens, which are used to train “M2”, a language model that directly reads and writes neural-compressed text.
Given a perfect probabilistic model of the raw byte sequence, the compression step would output a fully-compressed bitstream that would be indistinguishable from random noise, and hence unlearnable by M2. In reality, M1 can never be perfect (Zvonkin & Levin, 2007), so the M1-compressed output will still contain learnable patterns. We explore whether using compression powered by a relatively small M1 is able to “remove” the simple structure that M1 understands from the input—e.g., patterns of spelling, word frequency, and basic grammar—while retaining any higher-level structure that M1 fails to model—e.g., patterns requiring “deeper” reasoning and long range coherence. A larger M2 would then learn to model this higher-level structure, without needing to relearn the low-level structure removed by M1.444Intuitively, training M2 could be seen as analogous to fitting the residuals of M1 (Friedman, 2001). In theory, this process could be repeated by training an even-larger M3 model on text compressed by M2, and so on.
@@
At the same time, our best M2 models underperform subword baselines. We suspect this is due at least in part to the relatively unstable mappings our neural tokenizers induce between words and tokens. By contrast, standard subword tokenizers induce essentially stable word-to-token mappings, which likely makes the token sequences they output well-suited for LLM training. We illustrate this contrast through qualitative examples. Whether a neural tokenizer can reach a high level of compression while maintaining high learnability for LLM training is an interesting question for future research.
Our main contributions are as follows: (1) Outline advantages and challenges of training over neurally compressed text. (2) Compare LLMs trained over different tokenizers along two axes: bits/byte and FLOPs/byte. (3) Show that standard LLMs can’t learn to model vanilla AC-compressed text. (4) Show that GZip-compressed text is learnable by standard LLMs, but not competitive. (5) Propose compression via Equal-Info Windows, and show that it enables learning over neurally compressed text.
@@ 2 Motivation and Background
2.1 Advantages of Training over Neurally Compressed Text
@@
The most straightforward advantage is efficiency. By compressing the same text into a shorter token sequence, the model can process more text for the same computational cost. In particular, a model trained over $C{\mkern-1.0mu\times\mkern-1.0mu}{}$ compressed text will see $C{\mkern-1.0mu\times\mkern-1.0mu}{}$ more text during training compared to a model trained over raw text, given an equal compute budget. Increasing the amount of data seen in pretraining is often an effective means of improving performance (Kaplan et al., 2020, Hoffmann et al., 2022). Processing text more efficiently also confers benefits at inference time, reducing the serving cost for handling a request of a given prompt and continuation length. In addition to reducing the raw compute needed for inference, compression can also improve inference latency, since generating better-compressed output requires fewer sequential autoregressive steps.
@@
A second advantage is that working with compressed text allows modeling longer contextual dependencies. In vanilla transformer-based models, computation for the self-attention layer scales quadratically with the sequence length, $O(n^{2}d)$. This has limited the sequence lengths used by such models in practical settings to ~$10$k tokens. [^5] Exploring sub-quadratic attention mechanisms is an area of active research (Ainslie et al., 2020, Wang et al., 2020, Kitaev et al., 2020, Zaheer et al., 2020, Beltagy et al., 2020, Child et al., 2019, et alia). However, regardless of the cost of attention, compressing the input increases the effective context “for free”. If, via compression, each token represents (on average) $C$ bytes of raw text, then the resulting LLM can model dependencies across $C{\mkern-1.0mu\times\mkern-1.0mu}{}$ longer distances compared to a raw text model operating over the same token sequence length. While the benefits of modeling longer context (beyond ~$1{,}000$ bytes) are modest when viewed merely as perplexity gains (Press et al., 2022), the ability to condition on long context is critical for many applications, such as retrieving content from a document, or answering a coding question provided documentation.
@@
A third potential advantage of training over compressed text is that information will be spread more uniformly across the sequence. By the nature of compression, a text span that is relatively predictable (e.g., a boilerplate notice) will be more compressible than a span with high perplexity (e.g., a unique product serial number). When an LLM is trained over well-compressed text, each token will represent roughly an equal amount of information. Since the LLM allocates equal compute to each token, this amounts to allocating more compute for “harder” text spans. This adaptivity is similar in spirit to “Adaptive Computation Time” (ACT) (Graves, 2017), which learns to allocate additional compute at some sequence positions in an end-to-end manner, but with the advantage that in our case the computation remains “dense”—identical operations are applied at each position. [^6] It should be noted that ACT learns to allocate more compute where it is useful, as opposed to merely where the predictions are hard. For example, ACT learns to not waste compute on inherently unpredictable text spans. We expect that as a heuristic, allocating more compute to higher-perplexity text spans is valuable, but leave this to future work to verify.
@@ 2.2 Challenges of Training over Compressed Text
As a prerequisite for M2 to effectively predict continuations of compressed text, we anticipate that it is necessary for M2 to have the ability to decompress bits $\rightarrow$ text and compress text $\rightarrow$ bits. These sub-tasks are challenging in their own right. First, M2 needs to accurately “simulate” M1 in order to know the probabilities it assigns to the text, which determine the output of compression. [^7] For Arithmetic Coding, not only would M2 need to know the probabilities M1 assigns to the observed text, but it would also need to know the probabilities assigned to many unobserved symbols. This is because Arithmetic Coding operates over cumulative probabilities, i.e., the probability that the next symbol is e or any alphabetically preceding symbol. Training models to mimic other models can be difficult (Lester et al., 2022), and even in settings where models do learn to copy the behavior of another network (Hinton et al., 2015), this is often only when looking at which symbol was assigned the highest probability—the actual probabilities assigned often differ (Stanton et al., 2021). Second, M2 needs to learn the compression procedure itself. In our case, this means tracking the Arithmetic Coding algorithm, which requires maintaining high-precision numerical state across long contexts. We investigate these sub-tasks in detail in Section 5.2.
A further learnability challenge is the high level of context sensitivity needed to interpret a bitstream of compressed text. When chunked into tokens, a particular bit subsequence (e.g., 10111001) can map onto the same token despite having no stable “meaning” across occurrences. We show examples in Section 6.1, where a token maps to many different underlying text forms, necessitating strong contextual understanding. While LLMs are robust to some level of polysemy, as highlighted by the success of Hash Embeddings (Tito Svenstrup et al., 2017) where multiple unrelated words share a single token representation, we suspect this has its limits.
@@
Finally, a specific challenge of training over neurally compressed text is that multiple models need to be stored and run side-by-side in order to perform inference. We assume that if M1 is relatively small, this additional overhead is not a significant drawback compared to a standard tokenizer, which is also a separate model that is needed to tokenize text input and detokenize LLM outputs. In evaluating our approach, we include M1 compute in our calculations of total inference cost (FLOPs/byte).
@@ 2.3 Compression
In this work, we focus on lossless compression, which aims to encode a sequence of input symbols, $x_{0:N}=\{x_{0},x_{1},\dots,x_{N}\}\in X^{|V|}$, into a bitstream while minimizing the expected length of the bitstream. Compression methods are often factored into a “modeling” component and a “coding” component (Mahoney, 2013). The input sequence can be viewed as a sample from a true distribution $p$, $x_{0:N}\sim\leavevmode\nobreak\ p$, with a standard autoregressive decomposition, $p(x_{0:N})=\prod_{i=1}^{N}p(x_{i}|x_{0},\dots,x_{i-1})$. The “modeling” component aims to approximate $p$ with $\hat{p}$. While some compression algorithms assume static probabilities for each symbol, stronger algorithms are “adaptive”, meaning that symbol probabilities may change based on context. In this work, we use context-aware transformer-based language models to represent $\hat{p}$.
The “coding” component of a compression algorithm converts the input sequence to a bitstream of length $\ell(x_{0:N})$. To maximize compression, we want a coding algorithm that minimizes the expected number of bits in the bitstream, $L\coloneqq\mathbb{E}_{x_{0:N}\sim p}[\ell(x_{0:N})]$. This is done by assigning shorter bit sequences to common symbols and longer sequences to less common ones. [^8] This process can result in extremely uncommon sequences becoming longer under compression, as no algorithm can compress all possible input strings (Mahoney, 2013). In practice, natural language inputs are highly compressible and these edge cases are inputs that one would not recognize as natural language. The expected length is lower bounded by $L\geq H(p)$ where $H(p)\coloneqq\mathbb{E}_{x_{0:N}\sim p}[-\log_{2}p(x)]$ (Shannon, 1948). This means that, given a near-optimal coding algorithm, the achievable level of compression derives from how well the model $\hat{p}$ approximates $p$.
@@ 2.4 Arithmetic Coding
Arithmetic Coding (Rissanen, 1976, Pasco, 1977) uses a model $\hat{p}$ to compresses a sequence $x_{0:N}$ to a bitstream, which is the binary expansion of a float $f\in[0,1)$. The float $f$ is found by assigning successively smaller sub-intervals to each symbol $x_{i}\in x_{0:N}$, with the final interval enclosing $f$. An interval is made of an upper and lower bound, $I_{i}=[l_{i},u_{i})$ and its size is given by $u_{i}-l_{i}$. Starting with $I_{0}=[0,1)$, at each step of encoding, the interval for the symbol $x_{i}$ is created by partitioning the interval $I_{i-1}$ based on the cumulative distribution of $\hat{p}$ given the previous context, $\hat{p}_{cdf}(x_{i}|x_{<i})$. The size of this interval is given by $\text{size}(I_{i-1})*\hat{p}(x_{i}|x_{<i})$. Thus:
@@
The coding component of Arithmetic Coding is nearly optimal: the output bitstream will have a length of $-\lceil\log\hat{p}(x_{0:N})\rceil+1$ bits when using infinite precision. In the finite precision setting using $\beta$ bits, an extra $O(N2^{-\beta})$ bits are added (Howard & Vitter, 1992). See (Witten et al., 1987) for an example implementation. In our experiments, we use precision $\beta=14$. The practical effect of using a finite precision implementation of Arithmetic Coding is that the model’s cumulative distribution gets quantized to integers using $\beta$ bits. This results in a minimum probability of $2^{-\beta}$ being assigned to all tokens.
@@ 2.5 Related Work
Recent work has looked at using large language models for compression, but has not to our knowledge attempted to train subsequent models over the resulting compressed output. Works like (Delétang et al., 2024) use a transformer language model as the modeling component of Arithmetic Coding, but they do not train over compressed output nor do they make modifications to the compression algorithm to facilitate learnability by downstream models. Additionally, they focus on the setting of compressing fixed-size sequences of bytes. By contrast, our models operate over input sequences of fixed token length. This allows for models with higher compression rates to leverage longer contexts, as more bytes are included in the input.
@@
(Valmeekam et al., 2023) proposes changes to Arithmetic Coding to make it more amenable to use with LLMs—namely, they rank sort the logits from the model before creating text intervals, $I_{i}(x_{0:N})$. This could help alleviate issues stemming from errors in M2’s simulation of M1. However, they do not train models on top of their compressed output.
@@
Some approaches to “token-free” (i.e., purely character- or byte-level) language modeling down-sample the input sequence via convolutions (Clark et al., 2022, Tay et al., 2022), which could be seen as a form of end-to-end neural tokenization. However one important distinction is that the resulting tokenization is “soft”—outputting high-dimensional vectors and not implying a discrete segmentation—in contrast to our tokenization that outputs discrete tokens.
@@
Methods for learning discrete tokenization end-to-end have also been proposed (Chung et al., 2017, Godey et al., 2022). In the case of MANTa (Godey et al., 2022), the learned segmentation appears to be fairly semantic (i.e., respecting word and morpheme boundaries), which could be an advantage over our approach. However, they lack our bias towards encoding an equal amount of information per token.
@@
In modeling audio, it is common practice to use learned tokenizers that compress the raw input signal to discrete tokens from a fixed-size codebook (van den Oord et al., 2017, Baevski et al., 2020, Chung et al., 2021, Borsos et al., 2023). However, this compression is lossy, whereas we focus on lossless compression.
@@
Other recent work focuses on using the “modeling” component from well-known compressors to do other tasks. (Jiang et al., 2022) uses the model from GZip to perform text classification. (Vilnis et al., 2023) uses the Arithmetic Decoding algorithm with an LLM as the model to do diverse parallel sampling from that LLM. One could imagine that the “model” of our compressors (M1) is a teacher for M2, but unlike these other applications, the M1 values are not used outside of compression.
@@
(Külekci, 2011) also explores learning over compressed text, but with several key differences. First, they use n-gram language models (Shannon, 1948) while we use LLMs. Second, their model is conditioned on compressed bitstreams but produces a distribution over the raw, uncompressed, bytes while our M2 models predict directly in the compressed space. Additionally, they only consider static Huffman coding (Huffman, 1952) as the algorithm to compress model inputs. While this avoids the context sensitivity issues we outline in Section 2.2, it results in a far worse compression rate compared to the adaptive compression methods we use. One important distinction is that their equal-information windows are overlapping, and used as a sliding window to provide context to their n-gram language model. By contrast our equal-information windows are non-overlapping, and used to segment text into a series of equal-length bitstrings that can be interpreted independently by M2, and whose boundaries are easily identifiable, as they map to a fixed number of M2 tokens.
@@
Concurrently, (Goldman et al., 2024) explores how the compression performance of a tokenizer correlates with downstream model performance. They find that tokenizers that compress better perform better, which generally aligns with our findings, particularly in the large vocabulary setting, see Fig. 6. However, we find that using the strongest compressors is detrimental to learnability, as seen in the AC line in Fig. 3. These conflicting results likely stem from differences in tokenization strategy. Their work is restricted to BPE-based compressors while we explore stronger compressors built on LLMs and Arithmetic Coding. The qualitative differences between these classes of tokenizers are explored more in Section 6.1.